. Find the term independent of x in the expansion of (x^ 3+2/x)^ 20


1
Expert's answer
2021-06-21T07:58:30-0400

By the Binomial Theorem


"(a+b)^n=\dbinom{n}{0}a^nb^0+\dbinom{n}{1}a^{n-1}b^{1}+\dbinom{n}{2}a^{n-2}b^{2}"

"+...+\dbinom{n}{n-1}a^{1}b^{n-1}+\dbinom{n}{n}a^{0}b^{n}"

We have "a=x^3, b=\dfrac{2}{x}=2x^{-1}, n=20."

If the term is independent of "x," then


"(x^3)^{20-k}(x^{-1})^k=x^{0},"

"60-3k-k=0"

"k=15"

"\dbinom{20}{15}(x^3)^{20-15}\bigg(\dfrac{2}{x}\bigg)^{15}=15504(32768)"

"=508035072"